# MapNav: A Novel Memory Representation via Annotated Semantic Maps for VLM-based Vision-and-Language Navigation

### v2

@@ Abstract
Lingfeng Zhang1,2,thanks:  Co-first Authors.      Project Leader.  🖂 Corresponding Author.         Xiaoshuai Hao2,∗,†        Qinwen Xu2,5         Qiang Zhang1,3        Xinyao Zhang1 Pengwei Wang2      Jing Zhang4      Zhongyuan Wang2      Shanghang Zhang🖂${}^{2,5,}\textsuperscript{\Letter}$      Renjing Xu1,🖂 1The Hong Kong University of Science and Technology (Guangzhou) 2 Beijing Academy of Artificial Intelligence 3Beijing Innovation Center of Humanoid Robotics Co., Ltd. 4 School of Computer Science, Wuhan University 5 State Key Laboratory of Multimedia Information Processing, School of Computer Science, Peking University
@@ 1 Introduction
Vision-and-language Navigation (VLN) Gu et al. (2022); Park and Kim (2023) is a crucial integration of embodied AI and multimodal understanding, empowering autonomous agents to interpret natural language instructions and navigate complex, unseen environments. Unlike traditional navigation tasks that rely solely on visual input or predetermined waypoints, VLN requires a complex blend of language understanding and visual perception. For example, agents must interpret subtle instructions Vasudevan et al. (2021); An et al. (2024); Chen et al. (2024) such as “walk past the potted plants and turn left in front of the wooden cabinet”, while also processing dynamic visual scenes and making real-time navigation decisions.
@@ 2 Related Work
Vision-and-Language Navigation Vision-and-Language Navigation (VLN) Gu et al. (2022); Park and Kim (2023) in embodied AI focuses on navigating unseen environments by following human instructions, primarily in discretized simulated scenarios Ku et al. (2020b); Thomason et al. (2020). Agents navigate between predefined nodes on a graph by integrating language and visual input Qi et al. (2020); Liu et al. (2024), but this reliance complicates real-world deployment. To address this, VLNs for Continuous Environments (VLN-CE) Krantz et al. (2020); Savva et al. (2019) enable unrestricted navigation through low-level control or waypoint-based methods Hong et al. (2022); Krantz et al. (2021), improving sim-to-real transferability despite added complexity. Recent advancements in vision-language models have significantly influenced VLN development, utilizing large-scale pre-trained models Zhang et al. (2024a); Zheng et al. (2024) and VLN-specific pre-training Hao et al. (2020); Wu et al. (2022). For instance, NavGPT Zhou et al. (2024) autonomously generates actions using GPT-4o, while DiscussNav Long et al. (2024b) employs VLN experts to reduce human involvement. InstructNav Long et al. (2024a) decomposes navigation into subtasks, and Nav-CoT Lin et al. (2024) uses Chain of Thoughts (CoT) Wei et al. (2022) for environmental simulation. Some methods Zheng et al. (2024); Zhang et al. (2024a) fine-tune VLMs for specific navigation tasks, highlighting flexibility. However, existing approaches often depend on hierarchical prompts or historical frames, leading to high memory demands and limited understanding of past data. This paper introduces a novel memory representation using Annotated Semantic Maps (ASMs) to effectively replace traditional historical frames.
Map Representations for VLN Structured maps in VLN enhance navigation performance by improving environmental understanding Wang et al. (2023); Hong et al. (2023). Methods like MC-GPT Zhan et al. (2024) and VoroNav Wu et al. (2024) utilize topological maps to capture viewpoints and spatial relationships, while InstructNav Long et al. (2024a) and VLFM Yokoyama et al. (2024) create value maps for waypoint selection. Semantic maps Zhang et al. (2024c, b); Hong et al. (2023); Yu et al. (2023) retain object-level information for navigation. However, these maps are often not interpretable by VLMs. To address this, we propose Annotated Semantic Map (ASM), a novel semantic map representation that allows VLMs to explicitly understand rich map information, including obstacle distributions, explored areas, agent positions, historical trajectories, and semantic object locations. ASM aims to establish a new memory representation for VLN.
@@ 3.2 Our MapNav Agent
Inputs Encoding. MapNav utilizes an advanced vision-language architecture based on LLaVA-Onevision Li et al. (2024). At each timestep $t$, we process two primary visual inputs: the observation frame $\mathbf{X}_{t}$ and the ASM $\mathbf{M}_{t}$. These inputs are concurrently handled by a shared vision encoder, followed by modality-specific projectors for optimal feature alignment. Employing a SigLIP Zhai et al. (2023) vision encoder, we derive initial visual embeddings for both inputs. The observation frame produces $\mathbf{X}_{t}\in\mathbb{R}^{N\times D}$, where $N$ represents the number of image patches and $D$ is the model’s hidden dimension. Similarly, the ASM generates $\mathbf{X}^{M}_{t}\in\mathbb{R}^{N\times D}$. Both visual features undergo parallel spatial-aware transformations through a patch merge operation:
@@ 4.1 Experimental Details
Implementation Details. Our model is trained on a cluster server with 8 NVIDIA A100 GPUs for about 30 hours, totaling 240 GPU hours. We employ the LLaVA-Onevision Li et al. (2024) architecture, using Google’s SigLIP-so400M Zhai et al. (2023) as the vision encoder, Qwen2-7B-Instruct Wang et al. (2024) as the language model backbone, and Mask2Former Cheng et al. (2022) for semantic segmentation during ASM generation. For detailed experimental details, see Appendix B.
@@ 4.3 Ablation Studies
Efficiency Analysis. For the efficiency analysis, we compare memory consumption and inference time between MapNav and Navid. As shown in Tab. 3, MapNav demonstrates significant improvements in both areas. Specifically, our semantic map-based method maintains a constant memory footprint of 0.015MB, regardless of trajectory length, while Navid’s frame-based approach scales linearly, reaching 276MB at 300 steps. This is because MapNav only stores and updates a compact ASM, whereas Navid accumulates all historical RGB observations. In terms of inference speed, MapNav reduces average processing time by 79.5% (from 1.22 seconds per step to 0.25 seconds). This improvement arises because MapNav only calculates features from the current RGB frame and ASM, avoiding the need to process all historical frames.
@@ 5 Conclusion
In this paper, we introduce MapNav, a novel method using Annotated Semantic Maps (ASMs) for Vision-and-Language Navigation. It achieves state-of-the-art performance while significantly reducing memory usage compared to traditional approaches. By converting RGB-D and pose data into enriched top-down maps with annotations, MapNav integrates efficiently with Vision-Language Models. Experiments show it delivers superior navigation while maintaining a constant memory footprint of 0.015MB, regardless of trajectory length.
@@ Appendix A Dataset Construction
Incorporating General Vision-Language Datasets. To enhance our model’s visual understanding and general reasoning capabilities, we implement a co-training strategy that leverages both navigation-specific data and general-purpose vision-language datasets. Following the approach similar to LLaVA-OneVision Li et al. (2024), we incorporate a diverse set of high-quality visual understanding datasets alongside our navigation training data. Specifically, for R2R task, the co-training process utilizes approximately 600k samples from general vision-language tasks, which complement our 500k navigation training samples. The auxiliary dataset encompasses diverse visual understanding tasks including Visual Question Answering (VQA) samples for enhanced visual reasoning, multi-image reasoning tasks for improved cross-image understanding and relationship inference, and video-based tasks for strengthening temporal reasoning and dynamic scene comprehension. For more details, please refer to Li et al. (2024).
@@ Appendix B Details of Training
Model Setting. Our model builds upon the LLaVA-Onevision Li et al. (2024) framework and consists of three main components: the visual encoder, projector, and large language model (LLM). For the visual encoder, we employed Google’s SigLIP-so400m-patch14-384 Zhai et al. (2023), which processes input images at a resolution of 384x384 using patches of size 14x14. The SigLIP model serves as our vision backbone, converting input images into visual embeddings with a hidden dimension of 1152. The projector is implemented as a two-layer MLP with GELU activation, which maps the visual features to the language model’s embedding space. For the language model, we utilized Qwen2-7B-Instruct Wang et al. (2024) as our backbone, which features 28 transformer layers with 28 attention heads and a hidden size of 3584. The model supports a context length of up to 32,768 tokens and incorporates sliding window attention with a window size of 131,072 tokens.
@@ Appendix E More Experiments
Performance on Object Goal Navigation. To evaluate MapNav’s capability in handling sparse instructions, we conduct experiments on the zero-shot object goal navigation task using the Habitat-Matterport 3D (HM3D) Ramakrishnan et al. (2021) dataset. For this task, we modify the instruction format to: "Search for [object], move close to the object, and stop." Success requires the agent to achieve a Euclidean distance of no more than 1 meter from any instance of the target object category while maintaining visual line of sight. As shown in Tab. 7, despite not being trained on either the HM3D dataset or this instruction format, MapNav demonstrates strong generalization capabilities, achieving 34.6% SR and 25.6% SPL. This significantly outperforms WS-MGMap and Navid, which also utilize historical information and VLN training data.
@@
Table 7: Comparison of object goal navigation task under zero-shot setting. The best and second-best results are highlighted in bold and underlined, respectively.
Method SR↑ SPL↑
WS-MGMap Chen et al. (2022) 13.1 10.4
ZSON Majumdar et al. (2022) 25.5 12.6
GoW Gadre et al. (2023) 32.0 18.1
ESC Zhou et al. (2023) 35.5 23.5
Navid Zhang et al. (2024a) 32.5 21.5
MapNav (Ours) 34.6 25.6
@@
Table 9: Comparison of different VLMs on 100 randomly sampled R2R Val-Unseen split, including closed-source models, open-source models’ zero-shot performance, and our MapNav’s performance.
VLN-CE R2R Val-Unseen
VLMs NE↓ OS↑ SR↑ SPL↑
GPT-4o OpenAI (2023) 12.5 10.2 6.5 3.9
Claude 3.5 Sonnet Anthropic (2024) 11.6 9.8 7.2 4.6
Qwen2-VL Wang et al. (2024) 15.8 6.0 3.2 1.3
MapNav (Ours) 5.6 45.6 34.9 32.8
@@
Evaluation on Different VLMs. We conducted a comparative study of various VLMs on 100 randomly sampled episodes from the R2R Val-Unseen split, as presented in Tab. 9. The results demonstrate a significant performance gap between general-purpose VLMs and our specialized MapNav approach. While state-of-the-art models like GPT-4o OpenAI (2023) and Claude 3.5 Sonnet Anthropic (2024) achieve moderate performance with Success Rates of 6.5% and 7.2% respectively, they fall considerably short of our MapNav’s performance. Qwen2-VL Wang et al. (2024) shows more limited capabilities in this specific navigation task with an SR of 3.2%. In contrast, our MapNav achieves substantially better results across all metrics, with a Navigation Error of 5.6m and Success Rate of 34.9%. This stark performance difference highlights that while general-purpose VLMs possess broad visual-language understanding capabilities, specialized architectures like MapNav are crucial for complex embodied tasks such as vision-and-language navigation.

### v3

@@ Abstract
Lingfeng Zhang1,2,3,thanks:  Co-first Authors.     lfzhang715@gmail.com, xshao@baai.ac.cn   Project Leader.  🖂 Corresponding Authors.  renjingxu@hkust-gz.edu.cn, shanghang@pku.edu.cn         Xiaoshuai Hao2,∗,†        Qinwen Xu2,6         Qiang Zhang1,4 Xinyao Zhang1       Pengwei Wang2      Jing Zhang5      Zhongyuan Wang2 Shanghang Zhang🖂${}^{2,6,}\textsuperscript{\Letter}$      Renjing Xu🖂${}^{1,}\textsuperscript{\Letter}$ 1The Hong Kong University of Science and Technology (Guangzhou) 2 Beijing Academy of Artificial Intelligence, 3 Tsinghua University 4Beijing Innovation Center of Humanoid Robotics Co., Ltd. 5 School of Computer Science, Wuhan University 6 State Key Laboratory of Multimedia Information Processing, School of Computer Science, Peking University
@@ 1 Introduction
Vision-and-language Navigation (VLN) Gu et al. (2022); Park and Kim (2023) is a crucial integration of embodied AI and multimodal understanding Hao et al. (2025b); Ji et al. (2025); Tang et al. (2025), empowering autonomous agents to interpret natural language instructions and navigate complex, unseen environments. Unlike traditional navigation tasks that rely solely on visual input or predetermined waypoints, VLN requires a complex blend of language understanding and visual perception. For example, agents must interpret subtle instructions Vasudevan et al. (2021); An et al. (2024); Chen et al. (2024) such as “walk past the potted plants and turn left in front of the wooden cabinet”, while also processing dynamic visual scenes and making real-time navigation decisions.
@@ 2 Related Work
Vision-and-Language Navigation Vision-and-Language Navigation (VLN) Gu et al. (2022); Park and Kim (2023) in embodied AI Li et al. (2024b); Zhang et al. (2025); Hao et al. (2025a) focuses on navigating unseen environments by following human instructions, primarily in discretized simulated scenarios Ku et al. (2020b); Thomason et al. (2020). Agents navigate between predefined nodes on a graph by integrating language and visual input Qi et al. (2020); Liu et al. (2024), but this reliance complicates real-world deployment. To address this, VLNs for Continuous Environments (VLN-CE) Krantz et al. (2020); Savva et al. (2019) enable unrestricted navigation through low-level control or waypoint-based methods Hong et al. (2022); Krantz et al. (2021), improving sim-to-real transferability despite added complexity. Recent advancements in vision-language models have significantly influenced VLN development, utilizing large-scale pre-trained models Zhang et al. (2024a); Zheng et al. (2024); Wu et al. (2025) and VLN-specific pre-training Hao et al. (2020); Wu et al. (2022). For instance, NavGPT Zhou et al. (2024) autonomously generates actions using GPT-4o, while DiscussNav Long et al. (2024b) employs VLN experts to reduce human involvement. InstructNav Long et al. (2024a) decomposes navigation into subtasks, and Nav-CoT Lin et al. (2024) uses Chain of Thoughts (CoT) Wei et al. (2022) for environmental simulation. Some methods Zheng et al. (2024); Zhang et al. (2024a) fine-tune VLMs for specific navigation tasks, highlighting flexibility. However, existing approaches often depend on hierarchical prompts or historical frames, leading to high memory demands and limited understanding of past data. This paper introduces a novel memory representation using Annotated Semantic Maps (ASMs) to effectively replace traditional historical frames.
Map Representations for VLN Structured maps in VLN enhance navigation performance by improving environmental understanding Wang et al. (2023); Hong et al. (2023); Hao et al. (2024a, 2025c, b). Methods like MC-GPT Zhan et al. (2024) and VoroNav Wu et al. (2024) utilize topological maps to capture viewpoints and spatial relationships, while InstructNav Long et al. (2024a) and VLFM Yokoyama et al. (2024) create value maps for waypoint selection. Semantic maps Zhang et al. (2024c, b); Hong et al. (2023); Yu et al. (2023) retain object-level information for navigation. However, these maps are often not interpretable by VLMs. To address this, we propose Annotated Semantic Map (ASM), a novel semantic map representation that allows VLMs to explicitly understand rich map information, including obstacle distributions, explored areas, agent positions, historical trajectories, and semantic object locations. ASM aims to establish a new memory representation for VLN.
@@ 3.2 Our MapNav Agent
Inputs Encoding. MapNav utilizes an advanced vision-language architecture based on LLaVA-Onevision Li et al. (2024a). At each timestep $t$, we process two primary visual inputs: the observation frame $\mathbf{X}_{t}$ and the ASM $\mathbf{M}_{t}$. These inputs are concurrently handled by a shared vision encoder, followed by modality-specific projectors for optimal feature alignment. Employing a SigLIP Zhai et al. (2023) vision encoder, we derive initial visual embeddings for both inputs. The observation frame produces $\mathbf{X}_{t}\in\mathbb{R}^{N\times D}$, where $N$ represents the number of image patches and $D$ is the model’s hidden dimension. Similarly, the ASM generates $\mathbf{X}^{M}_{t}\in\mathbb{R}^{N\times D}$. Both visual features undergo parallel spatial-aware transformations through a patch merge operation:
@@ 4.1 Experimental Details
Implementation Details. Our model is trained on a cluster server with 8 NVIDIA A100 GPUs for about 30 hours, totaling 240 GPU hours. We employ the LLaVA-Onevision Li et al. (2024a) architecture, using Google’s SigLIP-so400M Zhai et al. (2023) as the vision encoder, Qwen2-7B-Instruct Wang et al. (2024) as the language model backbone, and Mask2Former Cheng et al. (2022) for semantic segmentation during ASM generation. For detailed experimental details, see Appendix B.
@@ 4.3 Ablation Studies
Efficiency Analysis. For the efficiency analysis, we compare memory consumption and inference time between MapNav and Navid. As shown in Tab. 3, MapNav demonstrates significant improvements in both areas. Specifically, our semantic map-based method maintains a constant memory footprint of 0.17MB, regardless of trajectory length, while Navid’s frame-based approach scales linearly, reaching 276MB at 300 steps. This is because MapNav only stores and updates a compact ASM, whereas Navid accumulates all historical RGB observations. In terms of inference speed, MapNav reduces average processing time by 79.5% (from 1.22 seconds per step to 0.25 seconds). This improvement arises because MapNav only calculates features from the current RGB frame and ASM, avoiding the need to process all historical frames.
@@ 5 Conclusion
In this paper, we introduce MapNav, a novel method using Annotated Semantic Maps (ASMs) for Vision-and-Language Navigation. It achieves state-of-the-art performance while significantly reducing memory usage compared to traditional approaches. By converting RGB-D and pose data into enriched top-down maps with annotations, MapNav integrates efficiently with Vision-Language Models. Experiments show it delivers superior navigation while maintaining a constant memory footprint of 0.17MB, regardless of trajectory length.
@@
Acknowledgement
This work was supported by the National Science and Technology Major Project (No. 2022ZD0117800).
@@ Appendix A Dataset Construction
Incorporating General Vision-Language Datasets. To enhance our model’s visual understanding and general reasoning capabilities, we implement a co-training strategy that leverages both navigation-specific data and general-purpose vision-language datasets. Following the approach similar to LLaVA-OneVision Li et al. (2024a), we incorporate a diverse set of high-quality visual understanding datasets alongside our navigation training data. Specifically, for R2R task, the co-training process utilizes approximately 600k samples from general vision-language tasks, which complement our 500k navigation training samples. The auxiliary dataset encompasses diverse visual understanding tasks including Visual Question Answering (VQA) samples for enhanced visual reasoning, multi-image reasoning tasks for improved cross-image understanding and relationship inference, and video-based tasks for strengthening temporal reasoning and dynamic scene comprehension. For more details, please refer to Li et al. (2024a).
@@ Appendix B Details of Training
Model Setting. Our model builds upon the LLaVA-Onevision Li et al. (2024a) framework and consists of three main components: the visual encoder, projector, and large language model (LLM). For the visual encoder, we employed Google’s SigLIP-so400m-patch14-384 Zhai et al. (2023), which processes input images at a resolution of 384x384 using patches of size 14x14. The SigLIP model serves as our vision backbone, converting input images into visual embeddings with a hidden dimension of 1152. The projector is implemented as a two-layer MLP with GELU activation, which maps the visual features to the language model’s embedding space. For the language model, we utilized Qwen2-7B-Instruct Wang et al. (2024) as our backbone, which features 28 transformer layers with 28 attention heads and a hidden size of 3584. The model supports a context length of up to 32,768 tokens and incorporates sliding window attention with a window size of 131,072 tokens.